A paddleboat can travel 60 miles downstream in the same time it takes to travel 140 miles upstream. If the speed of the current is 10 miles per hour, what is the speed of the paddleboat in still water?
Answer by flame8855(424)

there is something wrong in the question it should be 60 miles upward and 140 downward

speed downward = v+c = v+10
speed upward = v-c=v-10
d1=v/t , t1= v1/d1
t2= v2/d2
t1=t2
v+10/140 = v-10/60
v=25mph